# VScan: Rethinking Visual Token Reduction for Efficient Large Vision-Language Models

### v2

@@ C License Information
  • GQA [25]: compositional visual question answering. This benchmark is released under the CC BY 4.0 license.
  • ScienceQA [44]: multimodal science questions with diagrams and text. This benchmark is released under the MIT license.
  • VQAv2 [23]: real-world image-based QA with balanced answers. This benchmark is released under the CC BY 4.0 license.
  • TextVQA [52]: reading and reasoning over scene text in images. This benchmark is released under the CC BY 4.0 license.
  • VizWiz [24]: real-world visual questions from blind users. This benchmark is released under the CC BY 4.0 license.
  • MMBench [43] / MMBench-CN [43]: multilingual multimodal reasoning. These benchmarks are released under Apache License 2.0.
  • MME [20]: fine-grained multimodal evaluation on object, OCR, and commonsense. This benchmark is released under Apache License 2.0.
  • POPE [35]: probing object hallucinations in vision-language models. This benchmark is released under the MIT license.
  • RefCOCO / RefCOCO+ [30], RefCOCOg [47]: referential expression grounding. These benchmarks are released under Apache License 2.0.
  • TGIF [27]: video QA with spatiotemporal reasoning. The license of this work is not specified.
  • MSVD [10]: short video captioning and QA. This benchmark is released under the MIT license.
  • MSRVTT [61]: large-scale video-text retrieval and QA. This benchmark is released under the MIT license.
  • ActivityNet-QA [64]: complex event-centric video QA. This benchmark is released under Apache License 2.0.
@@
  • LLaVA-1.5 [39] is released under the LLaMA 2 Community License.
  • LLaVA-NeXT [40] is released under Apache License 2.0.
  • Qwen-2.5-VL [6] is released under Apache License 2.0.
  • VideoLLaVA [37] is released under Apache License 2.0.
Code. Our codebase builds upon PyramidDrop [60], licensed under MIT, and FasterVLM [68], licensed under Apache 2.0.

### v3

@@ C License Information
  • GQA Hudson and Manning [2019]: compositional visual question answering. This benchmark is released under the CC BY 4.0 license.
  • ScienceQA Lu et al. [2022]: multimodal science questions with diagrams and text. This benchmark is released under the MIT license.
  • VQAv2 Goyal et al. [2017]: real-world image-based QA with balanced answers. This benchmark is released under the CC BY 4.0 license.
  • TextVQA Singh et al. [2019]: reading and reasoning over scene text in images. This benchmark is released under the CC BY 4.0 license.
  • VizWiz Gurari et al. [2018]: real-world visual questions from blind users. This benchmark is released under the CC BY 4.0 license.
  • MMBench Liu et al. [2024d] / MMBench-CN Liu et al. [2024d]: multilingual multimodal reasoning. These benchmarks are released under Apache License 2.0.
  • MME Fu et al. [2023]: fine-grained multimodal evaluation on object, OCR, and commonsense. This benchmark is released under Apache License 2.0.
  • POPE Li et al. [2023b]: probing object hallucinations in vision-language models. This benchmark is released under the MIT license.
  • RefCOCO / RefCOCO+ Kazemzadeh et al. [2014], RefCOCOg Mao et al. [2016]: referential expression grounding. These benchmarks are released under Apache License 2.0.
  • TGIF Jang et al. [2017]: video QA with spatiotemporal reasoning. The license of this work is not specified.
  • MSVD Chen and Dolan [2011]: short video captioning and QA. This benchmark is released under the MIT license.
  • MSRVTT Xu et al. [2016]: large-scale video-text retrieval and QA. This benchmark is released under the MIT license.
  • ActivityNet-QA Yu et al. [2019]: complex event-centric video QA. This benchmark is released under Apache License 2.0.
@@
  • LLaVA-1.5 Liu et al. [2024a] is released under the LLaMA 2 Community License.
  • LLaVA-NeXT Liu et al. [2024b] is released under Apache License 2.0.
  • Qwen-2.5-VL Bai et al. [2025] is released under Apache License 2.0.
  • VideoLLaVA Lin et al. [2024] is released under Apache License 2.0.
Code. Our codebase builds upon PyramidDrop Xing et al. [2025], licensed under MIT, and FasterVLM Zhang et al. [2024], licensed under Apache 2.0.